OpenStudy (anonymous):

polluted water flows into a pond. the concentration of pollutant,c, in the pond at time,t, minutes is modeled by the equation c(t)=9-90000(1/10000+3t). where c is measured in kilograms per cubic metre. a) when will the concentration of pollutant in the pond reach 6 kg/m^3 b) what will happen to the concentration of pollutant over time.

ganeshie8 (ganeshie8):

\[\large c(t) = 9-90000\left(\dfrac{1}{10000+3t}\right)\] like this ?